# A Survey of Large Language Model Empowered Agents for Recommendation and Search: Towards Next-Generation Information Retrieval

### v1

@@ 4.2.6. User Simulators
Usimagent (Zhang et al., 2024g) is a search user behavior emulator based on LLMs. USimAgent can simulate users’ query, click, and stop behaviors in search tasks, generating complete search sessions. Through empirical research on actual user behavior datasets, researchers found that USimAgent outperforms existing methods in generating queries and performs similarly to traditional methods in predicting user click and stop behaviors. Sekuli (Sekulić et al., 2024a) explores the use cases of user simulators in conversation search systems. User simulators that can automatically answer clarification questions in multiple rounds of conversations and evaluate the rationality of system search results. Additionally, user simulators can generate large amounts of training data, thereby enhancing the performance of conversational systems. BASES (Ren et al., 2024) is an innovative user simulation framework based on LLM agents, designed to simulate web search user behavior comprehensively. It is capable of generating unique user profiles on a large scale, leading to diverse search behaviors. Specifically, through learning from vast amounts of web knowledge, the large language model can achieve human-like intelligence and generalization capabilities. Chen et al. (Chen et al., 2024b) proposed an online shopping task named ChatShop, where an agent needs to interact with the shopper to understand their needs and preferences gradually. They used LLM agents simulated shoppers to test the capabilities of the agent and found that the simulated shopper makes the task more realistic, increasing its complexity and challenge like real human shoppers.
@@ 4.3. Benchmark and Datasets
Furuta et al. (Furuta et al., 2024) introduced a benchmark, CompWoB, to highlight the limitations of language model agents in sequential task compositions and to emphasize the need for robust and generalized LLM agents for task combinations. CompWoB is a benchmark consisting of 50 new composite web automation tasks that reflect more realistic scenarios. Experimental results show that existing language model agents (such as gpt-3.5-turbo or gpt-4) achieve an average success rate of 94.0% on basic tasks, but their success rate drops to 24.9% on composite tasks. Agentbench (Liu et al., 2023) includes diverse agent-task scenarios and assesses LLMs’ agent capabilities through task performance data. This benchmark not only pinpoints the strengths and weaknesses of LLMs in agent tasks but also provides a unified standard for comparison. WebArena (Zhou et al., 2023b) and MIND2WEB (Deng et al., 2024a) provide environments that replicate realistic web browsing experiences, incorporating fully functional websites from diverse domains, such as e-commerce and social forums, with URLs, open tabs, and keyboard and mouse interactions. Cocktail (Dai et al., 2024a) is a benchmark providing a comprehensive tool for evaluating information retrieval models in the LLM era. Cocktail comprises 16 different datasets covering various text retrieval tasks and domains, including a mix of human-written and LLM-generated corpora. The diversity of these datasets enables researchers to assess the performance of information retrieval models across different scenarios. To avoid potential biases from datasets previously included in LLM training, the authors introduce a new dataset named NQ-UTD. This dataset features queries originating from recent events, ensuring the fairness and validity of the evaluation.
Overall, current research has introduced various datasets and benchmarks to evaluate the performance of LLM agents in IR scenarios. Unlike conventional LLM evaluation datasets that primarily focus on textual information, these datasets often include real online browsing tasks and user interaction behaviors.
@@ 5.1. Recent Advance of Embodied Agents in Cyber Environment
In recent years, researchers have been exploring the next piece of the puzzle of realizing AGI. Reinforcement Learning (RL) and Embodied AI have emerged in response to the thriving research of LLM agents. Consequently, there has been a shift from ”Internet agents” that focus on learning from constant datasets collected from the Internet, towards ”Embodied agents” which enable LLM agents to learn through interactions with their surroundings (Duan et al., 2022). Just as its name implies, the term embodied grants the agent a physical substrate, manifesting in the real world as using robotic arms to measure and interact with the environment or in cyber environments as the ”driving software” of information systems. Web agents serve as one of the key representations of embodied agents within the cyber environment. These agents must not only learn to navigate websites by interacting with GUI (Graphical User Interface) functions but also respond effectively to feedback from both web systems and human users. Such advancement inherently involves information retrieval tasks, requiring agents to perform complex multi-step searching and multi-modal retrievals. As a result, embodied agents are evolving into a new testbed and paradigm for research in recommendation and search areas.
@@ 5.2. Why Embodied Agents Can Be Used for Recommendation and Search
Embodied Agent Designing for Recommendation and Search. Existing embodied agents within cyber environment primarily focused on interaction with GUI. These agents concentrate on GUI operations and task-solving on a wide range of information systems like websites, desktops, and smartphones, which are typical scenarios for recommendation and search tasks. This close relationship could inspire innovative approaches and solutions for the next generation of recommendations and search. WorkArena (Drouin et al., 2024) introduces the first environment that supports chat-based agent-user interactions for the development and evaluation of web agents. Through a chat interface, the real human user can exchange messages with web agents. This allows for information retrieval tasks where a specific answer is expected from the agent, but also more practical tasks where user instructions change over time. This advancement may inspire further research and discussion around recommendation and search tasks, as agents would need to complete information retrieval tasks and output specific actions or plans. SmartAgent (Zhang et al., 2024d) defines a new task known as embodied personalized learning, where embodied cognition and item recommendation are interlaced as in many real-world personal assistance scenarios. To address this, a novel reasoning paradigm called Chain-of-User-Thought (COUT) is proposed to align user feedback with embodied agent actions under the progressive thoughts from basic GUI navigation to explicit and implicit user requirements. Leveraging the COUT paradigm, SmartAgent demonstrates the first full-stage embodied personalized capabilities in collaboration with a new dataset called SmartSpot, which for the first time supports a range of diverse embodied personalized environments.
In summary, the combination of embodied agents with recommendation and search systems could lead to the creation of intelligent systems offering more personalized services, but may also realize truly bionic personal assistants. This could be a potentially important direction for the next generation of recommendation and search research.
@@ 5.3. Limitation and Promising Directions
Promising directions. These limitations also spur new research directions. One of them is leveraging the one-for-all capability of embodied agents for transferable recommendation or search. Currently, large-scale pre-trained recommendation system models require extremely high costs to transfer to new scenarios. In some new scenarios where the gap is relatively large, the current pre-training and fine-tuning (PEFT) paradigm finds it difficult to achieve better domain adaptation. The ability of embodied agents to understand dynamic changes in the environment may be able to help with this. Another promising direction is the use of lightweight embodied agents for on-device recommendation. LLMs are often too resource-intensive to deploy on mobile devices, hindering the implementation of personalized services. Embodied agents, with their smaller footprint, may provide a solution to this problem. Additionally, the demand for zero-shot services on end-user devices also poses new requirements for embodied agent capabilities, which could drive more research. Finally, as embodied agents extend the capabilities of LLMs to recommendation and search areas within personal settings, the issue of privacy preservation becomes more crucial. Balancing user behavior modeling and individual privacy protection while providing intelligent services will be a critical and urgent problem.
@@ 6. Open Problems and Future Directions
Hallucinations. Hallucinations in LLMs refer to situations where the content generated by LLMs is inconsistent with facts or cannot be verified. For LLM agents in the IR field, hallucinations can result in the retrieval of incorrect information, thereby degrading the user experience.. Hallucinations can be divided into factuality hallucinations and faithfulness hallucinations (Huang et al., 2023c). The former refers to the inconsistency between the content generated by the LLM and the verifiable facts of the real world, while the latter refers to the inconsistency between the content generated by the model and the user’s instructions. Researchers have proposed many methods to alleviate these two types of hallucinations. For factual hallucinations, solutions such as RAG (Gao et al., 2023a; Xu et al., 2024a; Ram et al., 2023), CoT (Wei et al., 2022) and consistency enhancing (Wang et al., 2022; Shi et al., 2023; O’Brien and Lewis, 2023) can constrain LLM to obtain data from given sources. For command-type hallucinations, prompt engineering, RLHF, and other alignment technologies (Sharma et al., 2023; Perez et al., 2022; Wei et al., 2023) can strengthen the consistency of execution.
@@
Deployment Cost. LLM agents are typically composed of multiple LLM modules, resulting in a higher number of invocations and model parameters compared to general LLM applications. So, deploying LLM agents requires substantial resources, resulting in longer inference latency, which imposes higher hardware requirements on the end devices and may impact users’ experience. Researchers have put a lot of effort into this issue; Wilkins et al. (Wilkins et al., 2024) optimize LLM efficiency by proposing a cost-based scheduling framework, while Avatar (Wu et al., 2025) and QueryAgent (Huang et al., 2024) utilize a tool or environmental feedback to reduce inference time.
@@
Multi-Modal Agent. Multi-modal LLM agents can integrate data from different modalities, such as images and text, to more comprehensively understand user queries. For instance, when querying about a product, a user might provide both a textual description and an image of the product. A multi-modal LLM agent can analyze both the text and image information simultaneously, better understanding the product’s features and thereby providing more accurate retrieval results (Xie et al., 2024; Zhang et al., 2024h). Recently, numerous studies have begun exploring how to leverage multi-modal capabilities to enhance LLM agents’ ability to operate on web pages (Zheng et al., 2024b; Furuta et al., 2023) and mobile devices (Nong et al., 2024; Wang et al., 2024c). These studies typically utilize visual perception capabilities to identify visual and textual elements of mobile applications accurately. This forms a foundational basis for the planning and executing of subsequent operation tasks.
@@
Domain Specific Agent. In specialized domains with substantial domain knowledge, general LLM agents often perform poorly due to their reliance on common knowledge and data sources. However, data in these specialized domains is relatively sparse, making it insufficient to fully train LLMs. Therefore, efficiently and economically transforming general LLM agents into domain-specific LLM agents is a crucial direction for future research. Currently, related research has already been applied to relatively niche but popular categories such as clinical (Van Veen et al., 2024; Li et al., 2023c; Russe et al., 2024; Yan et al., 2024; Jin et al., 2024b) and law legal (Sun et al., 2024; Li et al., 2024b). In the future, as LLM agents improve their ability to utilize sparse data, there may also be opportunities for improvement in fields that have struggled with effective information dissemination due to their niche nature, such as psychological counseling, job searching, etc.
@@
Multi Agent Interaction. In multi-agent systems, individual agents can solve complex problems through information sharing and collaboration. Different agents may possess distinct knowledge and skills, and by cooperating, they can leverage their strengths to enhance overall performance information retrieval (Choi and Yoo, 1998; Wondergem et al., 1998). Recently, studies (Park et al., 2023, 2024) explore the interaction and evolution processes among multiple LLM agents, and several multi-agent frameworks (Chen et al., 2023; Qian et al., 2023; Team, 2023) have been proposed. Infogent (Reddy et al., 2024) utilized multi-agent systems to simulate human cognition for gathering information from the web, while Chateval (Chan et al., 2023) and MAD (Liang et al., 2023) improve agents’ generating quality through multi-agent debate.
@@
Personalization. Different users often have varying needs and preferences when performing information retrieval. For instance, the search term ”Apple” might refer to Apple Inc. for a tech investor, whereas it might mean the fruit for an average person. By leveraging the memory module and long context processing capabilities of LLM agents, personalized information retrieval tailored to individual users can be achieved (Li et al., 2024c). One approach to do so is to treat the user’s interaction history as context and input it into the model (Mo et al., 2024; Richardson et al., 2023; Aknouche et al., 2012). Mo et al. (Mo et al., 2024) incorporate conversational and personalized elements with LLM agents to fulfill user needs through multi-turn interactions. They indicate that Personal Textual Knowledge Bases (PTKBs) can effectively enhance conversational information retrieval, allowing retrieval results to be more closely aligned with the user’s background.
@@ 7. Conclusions
In this paper, we take the pioneering step in reviewing the advanced application of LLM agents in recommender and search systems. We begin the survey with a brief introduction to the concepts of recommendation and search tasks, as well as LLM agents, offering newcomers a foundational overview and critical background knowledge. Then, we elucidate the intrinsic nature of LLM agents, detailing three specific aspects to underscore why LLM agents can be used for recommendation and search. Moreover, we provide a detailed introduction to how to apply LLM agents in each task. For each phase, we provide a detailed taxonomy to categorize the major techniques and roles, drawing connections among the existing publications. Additionally, we extensively discuss the potential of applying Embodied agents in both tasks, delving into the rationale behind the advanced development. Finally, we summarize several challenges and promising directions in this field, which are expected to guide potential future directions.

### v2

@@ 4.2.6. User Simulators
Usimagent (Zhang et al., 2024g) is a search user behavior emulator based on LLMs. USimAgent can simulate users’ query, click, and stop behaviors in search tasks, generating complete search sessions. Through empirical research on actual user behavior datasets, researchers found that USimAgent outperforms existing methods in generating queries and performs similarly to traditional methods in predicting user click and stop behaviors. Sekuli (Sekulić et al., 2024a) explores the use cases of user simulators in conversation search systems. User simulators that can automatically answer clarification questions in multiple rounds of conversations and evaluate the rationality of system search results. Additionally, user simulators can generate large amounts of training data, thereby enhancing the performance of conversational systems. BASES (Ren et al., 2024) is an innovative user simulation framework based on LLM agents, designed to simulate web search user behavior comprehensively. It is capable of generating unique user profiles on a large scale, leading to diverse search behaviors. Specifically, through learning from vast amounts of web knowledge, the LLM can achieve human-like intelligence and generalization capabilities. Chen et al. (Chen et al., 2024b) proposed an online shopping task named ChatShop, where an agent needs to interact with the shopper to understand their needs and preferences gradually. They used LLM agents simulated shoppers to test the capabilities of the agent and found that the simulated shopper made the task more realistic, increasing its complexity and challenge like real human shoppers.
@@ 4.3. Benchmark and Datasets
Furuta et al. (Furuta et al., 2024) introduced a benchmark, CompWoB, to highlight the limitations of language model agents in sequential task compositions and to emphasize the need for robust and generalized LLM agents for task combinations. CompWoB is a benchmark consisting of 50 new composite Web Automation tasks that reflect more realistic scenarios. Experimental results show that existing language model agents (such as gpt-3.5-turbo or gpt-4) achieve an average success rate of 94.0% on basic tasks, but their success rate drops to 24.9% on composite tasks. Agentbench (Liu et al., 2023) includes diverse agent-task scenarios and assesses LLM agent capabilities through task performance data. This benchmark not only pinpoints the strengths and weaknesses of LLMs in agent tasks but also provides a unified standard for comparison. WebArena (Zhou et al., 2023b) and MIND2WEB (Deng et al., 2024a) provide environments that replicate realistic web browsing experiences, incorporating fully functional websites from diverse domains, such as e-commerce and social forums, with URLs, open tabs, and keyboard and mouse interactions. Cocktail (Dai et al., 2024a) is a benchmark providing a comprehensive tool for evaluating information retrieval models in the LLM era. Cocktail comprises 16 different datasets covering various text retrieval tasks and domains, including a mix of human-written and LLM-generated corpora. The diversity of these datasets enables researchers to assess the performance of information retrieval models across different scenarios. To avoid potential biases from datasets previously included in LLM training, the authors introduce a new dataset named NQ-UTD. This dataset features queries originating from recent events, ensuring the fairness and validity of the evaluation.
Overall, researchers have introduced various datasets and benchmarks to evaluate the performance of LLM agents in IR scenarios. Unlike conventional LLM evaluation datasets that primarily focus on textual information, these datasets often include real online browsing tasks and user interaction behaviors.
@@ 5.1. Recent Advance of Embodied Agents in Cyber Environment
In recent years, researchers have been exploring the next piece of the puzzle of realizing AGI. Reinforcement Learning (RL) and Embodied AI have emerged in response to the thriving research of LLM agents. Consequently, there has been a shift from “Internet agents” that focus on learning from constant datasets collected from the Internet, towards “Embodied agents” which enable LLM agents to learn through interactions with their surroundings (Duan et al., 2022). Just as its name implies, the term embodied grants the agent a physical substrate, manifesting in the real world as using robotic arms to measure and interact with the environment or in cyber environments as the “driving software” of information systems. Web agents serve as one of the key representations of embodied agents within the cyber environment. These agents must not only learn to navigate websites by interacting with GUI (Graphical User Interface) functions but also respond effectively to feedback from both web systems and human users. Such advancement inherently involves information retrieval tasks, requiring agents to perform complex multi-step searching and multi-modal retrievals. As a result, embodied agents are evolving into a new testbed and paradigm for research in recommendation and search areas.
@@ 5.2. Why Embodied Agents Can Be Used for Recommendation and Search
Embodied Agent Designing for Recommendation and Search. Existing embodied agents within cyber environment primarily focused on interaction with GUI. These agents concentrate on GUI operations and task-solving on a wide range of information systems like websites, desktops, and smartphones, which are typical scenarios for recommendation and search tasks. This close relationship could inspire innovative approaches and solutions for the next generation of recommendation and search. WorkArena (Drouin et al., 2024) introduces the first environment that supports chat-based agent-user interactions for the development and evaluation of web agents. Through a chat interface, the real human user can exchange messages with web agents. This allows for information retrieval tasks where a specific answer is expected from the agent, but also more practical tasks where user instructions change over time. This advancement may inspire further research and discussion around recommendation and search tasks, as agents would need to complete information retrieval tasks and output specific actions or plans. SmartAgent (Zhang et al., 2024d) defines a new task known as embodied personalized learning, where embodied cognition and item recommendation are interlaced as in many real-world personal assistance scenarios. To address this, a novel reasoning paradigm called Chain-of-User-Thought (COUT) is proposed to align user feedback with embodied agent actions under the progressive thoughts from basic GUI navigation to explicit and implicit user requirements. Leveraging the COUT paradigm, SmartAgent demonstrates the first full-stage embodied personalized capabilities in collaboration with a new dataset called SmartSpot, which for the first time supports a range of diverse embodied personalized environments.
In summary, the combination of embodied agents with recommender and search systems could lead to the creation of intelligent systems offering more personalized services, but may also realize truly bionic personal assistants. This could be a potentially important direction for the next generation of recommendation and search research.
@@ 5.3. Limitation and Promising Directions
Promising directions. These limitations also spur new research directions. One of them is leveraging the one-for-all capability of embodied agents for transferable recommendation or search. Currently, large-scale pre-trained recommendation models require extremely high costs to transfer to new scenarios. In some new scenarios where the gap is relatively large, the current pre-training and fine-tuning (PEFT) paradigm finds it difficult to achieve better domain adaptation. The ability of embodied agents to understand dynamic changes in the environment may be able to help with this. Another promising direction is the use of lightweight embodied agents for on-device recommendation. LLMs are often too resource-intensive to deploy on mobile devices, hindering the implementation of personalized services. Embodied agents, with their smaller footprint, may provide a solution to this problem. Additionally, the demand for zero-shot services on end-user devices also poses new requirements for embodied agent capabilities, which could drive more research. Finally, as embodied agents extend the capabilities of LLMs to recommendation and search areas within personal settings, the issue of privacy preservation becomes more crucial. Balancing user behavior modeling and individual privacy protection while providing intelligent services will be a critical and urgent problem.
@@ 6. Open Problems and Future Directions
Hallucinations. Hallucinations in LLMs refer to situations where the content generated by LLMs is inconsistent with facts or cannot be verified. For LLM agents in the IR field, hallucinations can result in the retrieval of incorrect information, thereby degrading the user experience. Hallucinations can be divided into factuality hallucinations and faithfulness hallucinations (Huang et al., 2023c). The former refers to the inconsistency between the content generated by the LLM and the verifiable facts of the real world, while the latter refers to the inconsistency between the content generated by the model and the user’s instructions. Researchers have proposed many methods to alleviate these two types of hallucinations. For factual hallucinations, solutions such as RAG (Gao et al., 2023a; Xu et al., 2024a; Ram et al., 2023), CoT (Wei et al., 2022), and consistency enhancing (Wang et al., 2022; Shi et al., 2023; O’Brien and Lewis, 2023) can constrain LLM to obtain data from given sources. For command-type hallucinations, prompt engineering, RLHF, and other alignment technologies (Sharma et al., 2023; Perez et al., 2022; Wei et al., 2023) can strengthen the consistency of execution.
@@
Deployment Cost. LLM agents are typically composed of multiple LLM modules, resulting in a higher number of invocations and model parameters compared to general LLM applications. So, deploying LLM agents requires substantial resources, resulting in longer inference latency, which imposes higher hardware requirements on the end devices and may impact users’ experience. Researchers have put a lot of effort into this issue. Wilkins et al. (Wilkins et al., 2024) optimize LLM efficiency by proposing a cost-based scheduling framework, while Avatar (Wu et al., 2025) and QueryAgent (Huang et al., 2024) utilize tools or environmental feedback to reduce inference time.
@@
Multi-Modal Agent. Multi-modal LLM agents can integrate data from different modalities, such as images and text, to more comprehensively understand user queries. For instance, when querying about a product, a user might provide both a textual description and an image of the product. A multi-modal LLM agent can analyze both the text and image information simultaneously, better understand the product’s features, and thereby provide more accurate retrieval results (Xie et al., 2024; Zhang et al., 2024h). Recently, numerous studies have begun exploring how to leverage multi-modal capabilities to enhance LLM agents’ ability to operate on web pages (Zheng et al., 2024b; Furuta et al., 2023) and mobile devices (Nong et al., 2024; Wang et al., 2024c). These studies typically utilize visual perception capabilities to identify visual and textual elements of mobile applications accurately.
@@
Domain Specific Agent. In specialized domains with substantial domain knowledge, general LLM agents often perform poorly due to their reliance on common knowledge and data sources. However, data in these specialized domains is relatively sparse, making it insufficient to fully train LLMs. Therefore, efficiently and economically transforming general LLM agents into domain-specific LLM agents is a crucial direction for future research. Currently, related research has already been applied to relatively niche but popular categories such as clinical (Van Veen et al., 2024; Li et al., 2023c; Russe et al., 2024; Yan et al., 2024; Jin et al., 2024b) and legal (Sun et al., 2024; Li et al., 2024b). In the future, as LLM agents progressively improve their ability to utilize sparse data, there may be opportunities for enhancements in niche areas such as psychological counseling and job seeking.
@@
Multi Agent Interaction. In multi-agent systems, individual agents can solve complex problems through information sharing and collaboration. Different agents may possess distinct knowledge and skills, and by cooperating, they can leverage their strengths to enhance overall performance information retrieval (Choi and Yoo, 1998; Wondergem et al., 1998). Recently, studies (Park et al., 2023, 2024) have explored the interaction and evolution processes among multiple LLM agents, and several multi-agent frameworks (Chen et al., 2023; Qian et al., 2023; Team, 2023) have been proposed. Infogent (Reddy et al., 2024) utilized multi-agent systems to simulate human cognition for gathering information from the web. Chateval (Chan et al., 2023) and MAD (Liang et al., 2023) improve agents’ generating quality through multi-agent debate.
@@
Personalization. Different users often have varying needs and preferences when performing information retrieval. For instance, the search term “Apple” might refer to Apple Inc. for a tech investor, whereas it might mean the fruit for others. By leveraging the memory module and long context processing capabilities of LLM agents, personalized information retrieval tailored to individual users can be achieved (Li et al., 2024c). One approach to do so is to treat the user’s interaction history as context and input it into the model (Mo et al., 2024; Richardson et al., 2023; Aknouche et al., 2012). Mo et al. (Mo et al., 2024) incorporate conversational and personalized elements with LLM agents to fulfill user needs through multi-turn interactions. They indicate that Personal Textual Knowledge Bases (PTKBs) can effectively enhance conversational information retrieval, allowing retrieval results to align more closely with the user’s background.
@@ 7. Conclusions
In this paper, we take the pioneering step in reviewing the advanced applications of LLM agents in recommender and search systems. We begin the survey with a brief introduction to the concepts of recommendation and search tasks, as well as LLM agents, offering newcomers a foundational overview and critical background knowledge. Then, we elucidate the intrinsic nature of LLM agents, highlighting three specific aspects to underscore why LLM agents can be used for recommendation and search. Moreover, we provide a detailed introduction to how to apply LLM agents in each task. For each phase, we provide a detailed taxonomy to categorize the major techniques and roles, drawing connections among the existing publications. Additionally, we extensively discuss the potential of applying embodied agents in both tasks, delving into the rationale behind the advanced development. Finally, we summarize several challenges and promising directions in this field, which are expected to guide potential future directions.